OpenStudy (anonymous):

Solve the equation. –6 + 3x = –9 A. –1 B. –6 C. –5 D. –3

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

sorry i am not the greatest at math and never have been.... when helping can you please show me how not only just give the answer thank you

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

can you show me the steps please

OpenStudy (anonymous):

so i better understand and know how to do it

OpenStudy (anonymous):

-6 + 3x = -9 then add 6 to each side -6 + 3x = -9 +6 +6 3x = -3 divide 3 on each side 3x = -3 x = 1

OpenStudy (anonymous):

i mean x = -1

OpenStudy (anonymous):

okay thanks so much

OpenStudy (anonymous):

Any time :)

OpenStudy (anonymous):

A

OpenStudy (hana88):

a

OpenStudy (anonymous):

a

OpenStudy (anonymous):

- 1